Grade 12th pass9 grade science

When a bus driver travelling at a speed of 20m/s applied brakes and brings the bus to rest in 10 seconds, then retardation will be

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
DEAR STUDENT
 

Solution

 

In the given question 
u=20m/s
v=0m/s
t=10s
and a=?
we know the formula v =u + at
 so putting the values we get 
0 = 20 + a X 10
- 20 = 10 a
-2 = a 
we get acceleration is -2 so retardation will be +2 m/ sec sq.

THANKS